OpenStudy (anonymous):

How do you solve these 2 equations: u= 2x-3y v= -x+y

OpenStudy (anonymous):

hey @psymon! lol

OpenStudy (psymon):

Lol, heya there. Because of the variables.....this is a vector question? No idea, haha. Not enough context for me to have any idea what's going on o.o

OpenStudy (anonymous):

lol my bad ill post the actual entire question in a pic

OpenStudy (anonymous):

This is the question

OpenStudy (anonymous):

This is the solution from the solutions manual

OpenStudy (anonymous):

an this is how chegg decided to start off the problem differently

OpenStudy (psymon):

Gotcha. quite a bit ahead of what ive seen. I just bought a calc 3 text, so I bet I could look itup. Chegg?

OpenStudy (anonymous):

yeah chegg is like a homework help site lol

OpenStudy (anonymous):

how would u solve those 2 equations though

OpenStudy (psymon):

Ah, cool xD Well, I'll have to look it up, completely new to me. I just gotta find jacobian and stuff lik that in my textbook. Have you been to the lamar website? Professor who has all his class notes online?

OpenStudy (anonymous):

yeah but i forgot to look around :S lol. Also my test is tomorrow... this is the very last section im cramming through right now. Not too smart..

OpenStudy (psymon):

Yeah, not good. And the funky greek symbol im not used to. That's for partial derivative, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got through one of these problems already that is similar but i seem to be stuck on solving those 2 equations on this one idk why

OpenStudy (psymon):

Okay, yay, closer to making sense xD Alright, let's see what i can do with the first part.

OpenStudy (anonymous):

by looking at the solution they go from 2x-3y = u and -x+y=v (which is given) to -x = u+3v and y = v+x. I feel like its something simple im missing

OpenStudy (psymon):

Im thinking some sort of substitution. Im hopping back and forth, so Ill be a bit slow. But I do wanna take a look at it.

OpenStudy (anonymous):

np

OpenStudy (psymon):

Yeah, thats exactly what they did. they took -x + y = v and said y = v+x. Then they substituted that y value in for y in the second equation.

OpenStudy (psymon):

You kinda see how they got that? xD

OpenStudy (anonymous):

Haha yeah I just worked it out doing what you said and got the -x = u +3v. im still not sure why they left it as -x. Would I then plug the -x into the 1st equation to get y = v+x

OpenStudy (psymon):

yeah, you would have to do another substitution. Use x = y-v and sub that into the 2x equation. You'll getthe other value you need :3

OpenStudy (anonymous):

would u recommend using this substituion method or the chegg route

OpenStudy (psymon):

Substitution. More straightforward and probably more familiar to us, lol. I would have never tried anything like what Chegg did, although I can see why they did that.

OpenStudy (anonymous):

substitution it is..lol.. I understand it a lot better thanks to you.

OpenStudy (psymon):

Glad that part works xD I'll work my way over to the jacobian part slowly, lol.

OpenStudy (anonymous):

rofl ok. The determinant part is easy. b is eh im gonna try it now

OpenStudy (psymon):

Alright, Ill ignore that part then, unless its needed for part b xD Just havent been able to focus on it yet, lol.

OpenStudy (psymon):

Figure it out? :P

OpenStudy (anonymous):

Yes just finished. Thanks so much! on to the next one -.-. on the bright side i've got about 4-5 problems left and then i can get a couple hours rest before my test lol

OpenStudy (psymon):

Lol, okay. Yeah, I figured it out, just had to make sure if you did, lol. Wasn;t that bad it seems O.o

OpenStudy (psymon):

Just a bunch of subbing xD

OpenStudy (anonymous):

Lol yeah it was just the substitution that got me. I really shouldn't be struggling with that when I'm in cal 3.. .-.

OpenStudy (psymon):

Lol, hey, it happens. How was calc 1 and 2 for ya then?

OpenStudy (anonymous):

cal 2 was hell for me lol. Cal 1 wasn't too bad. and Cal 3 would be easier if it wasn't only a month long haha

OpenStudy (psymon):

Yeah, thats the brutal part x_x I can't wait til I take it, but it was a schedule conflict D: I still bought the textbook for it to study, though. I gotta start from the beginning, though, haha. At least the first part is easy it seems. Calc 2 really helped out my overall calc skills xD

OpenStudy (anonymous):

dude you'll be fine you've been helping me out with like everything lol

OpenStudy (psymon):

I just gotta study it first, haha. Atleast for my textbook, looks like youre at the end of chapter 14 and theres 16 chapters xD

OpenStudy (anonymous):

haha im using the 12th edition of thomas calculus. this is chapter 15.8. We're starting and finishing 16 in like 1 day next week, then final like 2 days after >.<

OpenStudy (psymon):

Oh xD So howd you get the solution manual for it?

OpenStudy (anonymous):

chegg has the solutions online and I also found a pdf with the solutions ;D

OpenStudy (psymon):

Wow, nice, haha. For me, the textbook I use has a website that does some minimal work for the odd numbered questions. Sometimes it helps a lot, sometimes not at all, but calculator + that works out. I just need to find out how to get my calculator to do more of the calc 3 stuff x_x

OpenStudy (anonymous):

i have a browser with bookmarks dedicated to online calculators O.o lol

OpenStudy (psymon):

All of those work for this stuff? xD

OpenStudy (anonymous):

not really rofl :(

OpenStudy (psymon):

Ah D: I think my calculator should be able to do some of it, I just need to learn how x-x

OpenStudy (anonymous):

@Psymon Got an 80 on my test! Wooooo